What percentage of the population live in their state of birth? According to the 2010 U.S. census, 14% of people born in Nevada, 26% of people born in Florida, 35% of people born in Virginia, 42% of people born in California, and 50% of people born in Arkansas live in that state. Suppose you wanted to see if these percentages still hold true today, so you randomly selected 40 people born in Florida and 12 said they live in that state. Based on your findings, can you conclude the proportion of all people born in Florida that live in that state is now lower at α=.005?

For the hypothesis stated above, what is the P-value?

a.

0.7088

b.

None of the answers is correct

c.

0.2912

d.

0.2810

e.

0.7190
Answer #1

Sample proportion = 12/40 = 0.30

This is a left tailed test.

Test statistic

z = 0.58

p - Value = P(z < 0.58) = 0.7190

Option E is correct.